the ratio of the volumes of three buckets is 3:4:5. Buckets contain the mixture of water and alcohol. If the mixture contains water and alcohol in the ratio 1:4 , 1:3, 2:5 respectively, then find the ratio of water and alcohol when the mixture in all containers is poured into a fourth container.
Answer by richwmiller(17219)

3/4*3x+4x/3+ 3/5*5x=12x*y
y = 79/144 alcohol to water
3x/4+4x/3+2/5*5x=12x*y
y = 49/144 water to alcohol